(1) Which of the following statements is correct for the equation shown here?

4NH3(g) + 5O2(g) ⟶ 4NO(g) + 6H2O(g)

(a) 6 g of H2O are produced for every 4 g of NH3 reacted. (b) 1 mole of NO is produced per mole of NH3 reacted. (c) 2 moles of NO are produced for every 3 moles of O2 reacted.

(2) Silicon reacts with chromium(III) oxide according to the equation

3Si(s) + 2Cr2O3(s) ⟶ 3SiO2(s) + 4Cr(s)

If 59.4 g of silicon is consumed in the reaction, what is the mass of SiO2 produced?
Answer #1

1)

Coefficient in stoichiometric equation represent ratio of mol or volume or number of particles. They do not represent ratio of mass.

So, a is wrong

From reaction,

Ratio of NH3 / NO = 4/4 = 1/1

So, b is correct

Ratio of O2/N2 = 5/4 = 1.25

In option c, ratio is 3/2 = 1.5

Hence this is not correct

Answer: b

2)

Molar mass of Si = 28.09 g/mol

mass of Si = 59.4 g

mol of Si = (mass)/(molar mass)

= 59.4/28.09

= 2.115 mol

According to balanced equation

mol of SiO2 formed = moles of Si

= 2.115 mol

Molar mass of SiO2,

MM = 1*MM(Si) + 2*MM(O)

= 1*28.09 + 2*16.0

= 60.09 g/mol

mass of SiO2 = number of mol * molar mass

= 2.115*60.09

= 127 g

Answer: 127 g